# Distributional Finite Element curl div Complexes and Application to Quad Curl Problems

### v4

@@ Abstract.
The paper addresses the challenge of constructing conforming finite element spaces for the ${\rm curl\,}\operatorname{div}$ operator in three dimensions. Tangential-normal continuity is introduced in order to develop distributional finite element ${\rm curl\,}\operatorname{div}$ complexes. The spaces constructed are applied to discretize the quad curl problem, demonstrating optimal order of convergence. Furthermore, a hybridization technique is proposed, demonstrating its equivalence to nonconforming finite elements and weak Galerkin methods.
@@
The first author was supported by NSF DMS-2012465, and DMS-2309785. The second author was supported by the National Natural Science Foundation of China Project 12171300, and the Natural Science Foundation of Shanghai 21ZR1480500.
@@ 1. Introduction
In this work, we will design tangential-normal continuous finite element spaces for the $2$nd order differential operator curldiv and construct the corresponding distributional finite element curldiv complexes in three dimensions. We will apply the constructed finite elements to solve the fourth-order curl problem $-{\rm curl\,}\Delta{\rm curl\,}\boldsymbol{u}=\boldsymbol{f},\operatorname{div}\boldsymbol{u}=0$ in a domain $\Omega\subset\mathbb{R}^{3}$ with boundary conditions $\boldsymbol{u}\times\boldsymbol{n}=\operatorname{curl}\boldsymbol{u}=0$ on $\partial\Omega$. Such problem arises from multiphysics simulation such as modeling a magnetized plasma in magnetohydrodynamics [5, 33, 10].
@@
Simple finite element spaces with tangential-normal continuity can be constructed. For an integer $k\geq 0$, take $\mathbb{P}_{k}(T;\mathbb{T})$ as the space of shape functions. The degrees of freedom (DoFs) are given by:
@@
where $\boldsymbol{t}_{1}$ and $\boldsymbol{t}_{2}$ denote two mutually perpendicular unit tangential vectors of face $F$ and are used to determine the tangential component of the vector $\boldsymbol{\tau}\boldsymbol{n}$. The global finite element space $\Sigma_{k,h}^{\rm tn}$ by requiring single valued (2a) is tangential-normally continuous. A different tangential-normally continuous finite element proposed in [22, 34] requires that $(\boldsymbol{t}_{i}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})|_{F}\in\mathbb{P}_{k-1}(F)$, meaning the tangential-normal trace is reduced by one degree. A bubble-enriched space, $\Sigma_{k,h}^{\rm tn}$, was introduced in [21] for the Stokes equation with weakly imposed stress symmetry. Our newly developed tangential-normal continuous finite element space $\Sigma_{k,h}^{\rm tn}$ can also be used to discretize a distributional formulation of the Stokes equation, which will be explored elsewhere.
@@
Furthermore, we apply hybridization techniques to (5a)-(5b) which is equivalent to a stabilization-free weak Galerkin method and to the $H({\rm grad\,}{\rm curl\,})$ nonconforming finite elements in [30, 48] for solving the quad curl problem. For other discretization of the quad-curl problem, we refer to the macro finite element method in [29], nonconforming finite element methods in [31, 42, 44], mixed finite element methods in [41, 11], decoupled finite element methods in [9, 47, 8], and references cited therein.
@@ 3. Distributional finite element ${\rm curl\,}\operatorname{div}$ complex
We shall construct a finite element counterpart of the distributional ${\rm curl\,}\operatorname{div}$ complex (8). The key is a tangential-normal continuous and traceless tensor finite element space.
@@ 3.1. Finite element spaces
Take $\mathbb{P}_{k}(T;\mathbb{T})$ as the space of shape functions with $k\geq 0$. The degrees of freedom (DoFs) are given by
@@
In order to prove the unisolvence, we first present two intrinsic bases of $\mathbb{T}$ which are variants of a basis constructed in [25].
@@
$\operatorname{tr}_{F}^{\rm tn}\boldsymbol{\tau}=\boldsymbol{n}_{F}\times\boldsymbol{\tau}\boldsymbol{n}_{F}$
and $\operatorname{tr}^{\rm tn}:C(T;\mathbb{T})\to L^{2}(\partial T;\mathbb{R}^{2})$ as $\operatorname{tr}^{\rm tn}|_{F}=\operatorname{tr}_{F}^{\rm tn}$.
Lemma 3.2.
The DoFs (11a)-(11b) are unisolvent for space $\mathbb{P}_{k}(T;\mathbb{T})$.
Proof.
The number of DoFs (11a)-(11b) is
$8{k+2\choose 2}+8{k+2\choose 3}=8{k+3\choose 3}=\dim\mathbb{P}_{k}(T;\mathbb{T}).$
Take a $\boldsymbol{\tau}\in\mathbb{P}_{k}(T;\mathbb{T})$, thanks to Lemma 3.1, we can express
$\boldsymbol{\tau}=\sum_{\ell=0}^{3}\big{(}c_{i\ell}(x)\operatorname{dev}(\nabla\lambda_{i}\otimes\boldsymbol{t}_{i\ell})+c_{j\ell}(x)\operatorname{dev}(\nabla\lambda_{j}\otimes\boldsymbol{t}_{j\ell})\big{)},$
where the coefficients $c_{i\ell}=\boldsymbol{t}_{mi}^{\intercal}\boldsymbol{\tau}\nabla\lambda_{\ell}$ and $c_{j\ell}=\boldsymbol{t}_{mj}^{\intercal}\boldsymbol{\tau}\nabla\lambda_{\ell}$ are polynomials in $\mathbb{P}_{k}(T)$.
Assume all the DoFs (11a)-(11b) vanish. As $\boldsymbol{t}_{1}^{\intercal}\boldsymbol{\tau}\boldsymbol{n}|_{F},\boldsymbol{t}_{2}^{\intercal}\boldsymbol{\tau}\boldsymbol{n}|_{F}\in\mathbb{P}_{k}(F)$, the vanishing DoF (11a) implies $\operatorname{tr}^{\rm tn}\boldsymbol{\tau}=0$.
By Lemma 3.1, we test with $\boldsymbol{t}_{mi}\otimes\nabla\lambda_{\ell}$ and restrict on face $F_{\ell}$ to single out the coefficient $c_{i\ell}$:
$c_{i\ell}|_{F_{\ell}}=\boldsymbol{t}_{mi}^{\intercal}\boldsymbol{\tau}\nabla\lambda_{\ell}|_{F_{\ell}}=0.$
Therefore $c_{i\ell}=\lambda_{\ell}q_{i\ell}$ for some $q_{i\ell}\in\mathbb{P}_{k-1}(T)$. Take $\boldsymbol{q}=q_{i\ell}\boldsymbol{t}_{mi}\otimes\nabla\lambda_{\ell}\in\mathbb{P}_{k-1}(T;\mathbb{T})$ in (11b) to conclude $q_{i\ell}=0$. As $i,\ell$ are arbitrary, we conclude $\boldsymbol{\tau}=0$. ∎
Define the bubble polynomial space of degree $k$
@@
Since the basis functions are expressed in terms of intrinsic geometric quantities—tangential and normal vectors along with barycentric coordinates—the defined finite element spaces are affine invariant, meaning the space remains unchanged under affine transformations. Thus, the traditional method of using a reference tetrahedron and transforming to the physical element can be applied, resulting in the same finite element space. In particular, curved elements can be handled using the Piola transform, as described in equations (3.76)-(3.77) of [36]. Extending the two-dimensional work of [4] to a distributional mixed finite element method specifically tailored for the quad-curl problem on domains with curved boundaries is an interesting topic for future research.
@@ 3.3. Characterization of null spaces
The first Poincaré inequality (21) can be found in [20, Lemma 3.4 and Theorem 3.6] and [36, Lemma 7.20]. Consequently $({\rm curl\,}\cdot,{\rm curl\,}\cdot)$ is an inner product on $K_{h}^{c}$ and the operator $({\rm curl\,}{\rm curl\,})_{h}$ is isomorphic.
@@ 4. Mixed finite element method of the quad-curl problem
where $\boldsymbol{f}\in H^{-1}(\operatorname{div},\Omega)\cap\ker(\operatorname{div})$ is known. Such problem arises from multiphysics simulation such as modeling a magnetized plasma in magnetohydrodynamics [5, 33, 10].
@@ 4.3. Error analysis
where $\widetilde{\boldsymbol{\sigma}}:={\rm grad\,}{\rm curl\,}\widetilde{\boldsymbol{u}}$. When $\Omega$ is convex, regularity (51) has been obtained in [30, Lemma A.1].

### v5

@@ Abstract.
The paper addresses the challenge of constructing finite element ${\rm curl\,}\operatorname{div}$ complexes in three dimensions. Tangential-normal continuity is introduced in order to develop distributional finite element ${\rm curl\,}\operatorname{div}$ complexes. The spaces constructed are applied to discretize the quad curl problem, demonstrating optimal order of convergence. Furthermore, a hybridization technique is proposed, demonstrating its equivalence to nonconforming finite elements and weak Galerkin methods.
@@
The first author was partially supported by NSF DMS-2012465, DMS-2309777, and DMS-2309785. The second author was supported by the National Natural Science Foundation of China Project 12171300, and the Natural Science Foundation of Shanghai 21ZR1480500.
@@ 1. Introduction
In this work, we will construct distributional finite element curldiv complexes in three dimensions, and apply it to solve the fourth-order curl problem $-{\rm curl\,}\Delta{\rm curl\,}\boldsymbol{u}=\boldsymbol{f},\operatorname{div}\boldsymbol{u}=0$ in a domain $\Omega\subset\mathbb{R}^{3}$ with boundary conditions $\boldsymbol{u}\times\boldsymbol{n}=\operatorname{curl}\boldsymbol{u}=0$ on $\partial\Omega$. Such problem arises from multiphysics simulation such as modeling a magnetized plasma in magnetohydrodynamics [31, 9].
@@
We will use the tangential-normal continuous finite element constructed in [20] for the discretization of $\Sigma^{\rm tn}$. For an integer $k\geq 0$, take $\mathbb{P}_{k}(T;\mathbb{T})$ as the space of shape functions. The degrees of freedom (DoFs) are given by:
@@
where $\boldsymbol{t}_{1}$ and $\boldsymbol{t}_{2}$ denote two mutually perpendicular unit tangential vectors of face $F$ and are used to determine the tangential component of the vector $\boldsymbol{\tau}\boldsymbol{n}$. The global finite element space $\Sigma_{k,h}^{\rm tn}$ by requiring single valued (2a) is tangential-normally continuous.
@@
Furthermore, we apply hybridization techniques to (5a)-(5b), leading to a stabilization-free weak Galerkin method and extending to the $H({\rm grad\,}{\rm curl\,})$ nonconforming finite elements introduced in [28, 45] for solving the quad-curl problem. Equivalently, we identify the complex that accommodates these nonconforming finite elements and generalize them to arbitrary orders.
For other discretization of the quad-curl problem, we refer to the macro finite element method in [27], nonconforming finite element methods in [29, 39, 41], mixed finite element methods in [38, 10], decoupled finite element methods in [8, 44, 7], and references cited therein.
@@ 3. Distributional finite element ${\rm curl\,}\operatorname{div}$ complex
We shall construct a finite element counterpart of the distributional ${\rm curl\,}\operatorname{div}$ complex (8).
@@ 3.1. Finite element spaces
We first recall the tangential-normal continuous finite element for traceless tensors in [20]. Take $\mathbb{P}_{k}(T;\mathbb{T})$ as the space of shape functions with $k\geq 0$. The degrees of freedom (DoFs) are given by
@@
In order to give a geometric decomposition of space $\mathbb{P}_{k}(T;\mathbb{T})$, we present two intrinsic bases of $\mathbb{T}$ which are variants of a basis constructed in [23].
@@
$\operatorname{tr}_{F}^{\rm tn}\boldsymbol{\tau}=\boldsymbol{n}_{F}\times\boldsymbol{\tau}\boldsymbol{n}_{F},$
and $\operatorname{tr}^{\rm tn}:C(T;\mathbb{T})\to L^{2}(\partial T;\mathbb{R}^{2})$ as $\operatorname{tr}^{\rm tn}|_{F}=\operatorname{tr}_{F}^{\rm tn}$. Let the bubble polynomial space of degree $k$
@@
Since the basis functions are expressed in terms of intrinsic geometric quantities—tangential and normal vectors along with barycentric coordinates—the defined finite element spaces are affine invariant, meaning the space remains unchanged under affine transformations. Thus, the traditional method of using a reference tetrahedron and transforming to the physical element can be applied, resulting in the same finite element space. In particular, curved elements can be handled using the Piola transform, as described in equations (3.76)-(3.77) of [33]. Extending the two-dimensional work of [4] to a distributional mixed finite element method specifically tailored for the quad-curl problem on domains with curved boundaries is an interesting topic for future research. ∎
@@ 3.3. Characterization of null spaces
The first Poincaré inequality (21) can be found in [19, Lemma 3.4 and Theorem 3.6] and [33, Lemma 7.20]. Consequently $({\rm curl\,}\cdot,{\rm curl\,}\cdot)$ is an inner product on $K_{h}^{c}$ and the operator $({\rm curl\,}{\rm curl\,})_{h}$ is isomorphic.
@@ 4. Mixed finite element method of the quad-curl problem
where $\boldsymbol{f}\in H^{-1}(\operatorname{div},\Omega)\cap\ker(\operatorname{div})$ is known. Such problem arises from multiphysics simulation such as modeling a magnetized plasma in magnetohydrodynamics [31, 9].
@@ 4.3. Error analysis
where $\widetilde{\boldsymbol{\sigma}}:={\rm grad\,}{\rm curl\,}\widetilde{\boldsymbol{u}}$. When $\Omega$ is convex, regularity (51) has been obtained in [28, Lemma A.1].